OpenStudy (anonymous):

A rectangular box is x feet long and x feet wide. The volume of the box is 4x8 + 3x6 cubic feet. What polynomial represents the height of the box in feet? Choices are: A. 4x + 3 B. 4x6 + 3x4 C. 4x8 + 3x6 D. 4x7 + 3x5

jimthompson5910 (jim_thompson5910):

Volume = Length * Width * Height V = LWH 4x^8 + 3x^6 = x*x*H 4x^8 + 3x^6 = x^2*H (4x^8 + 3x^6)/(x^2) = H H = (4x^8 + 3x^6)/(x^2) H = (x^2(4x^6 + 3x^4))/(x^2) H = (1(4x^6 + 3x^4))/(1) H = 1(4x^6 + 3x^4) H = 1*4x^6 + 1*3x^4 H = 4x^6 + 3x^4 There's a lot going on here, but the basic idea is to isolate H and simplify as much as you can.